2 Each of two small spheres is positively charged, the combined charge equal to +5.0x 103 C ne Coulomb force between these charges is 1.0 N when they are placed 2.0 meters apart. What is the value of each charge?
Answer #1

q1 + q2 = 5 *10^-5

q1 = 5*10^-5 - q2 (i)

also

F = k q1 q2 / r^2

1 = 9*10^9* q2 * ( 5 - q2)*10^-10 / 2^2

q2^2 - 5q2 + 4.444 = 0

q2 = 3.84*10^-5 C

putting in (i)

q1 = 1.16*10^-5 C

======

do comment in case any doubt, will reply for sure.. Goodluck